OpenStudy (anonymous):

The polynomial of degree 5, P(x) has leading coefficient 1, has roots of multiplicity 2 at x=3 and x=0, and a root of multiplicity 1 at x=-5

OpenStudy (campbell_st):

do you have to find the polynomial..?

OpenStudy (anonymous):

yes find the polynomial. I actaully figured this one out but can you help me with this. The polynomial of degree 4, P(x) has a root of multiplicity 2 at x=3 and roots of multiplicity 1 at x=0 and x=-4. It goes through the point (5,108). . how do i find a specific point?

OpenStudy (anonymous):

F(x)=(x-3)^2(x)(x+4)

OpenStudy (campbell_st):

yep thats good but seeing you don't know anything about the coefficient of the leading term you need \[y = a[x(x+4)(x-3)^2]\] substitute the values from the point to find the value of the constant a

OpenStudy (anonymous):

oh okay, that makes sense, thank you. The polynomial of degree 3, P(x), has a root of multiplicity 2 at x=4 and a root of multiplicity 1 at x=-5. The y-intercept is y= -56 . Find a formula for P(x) . do i plug the -56 in for the y?

OpenStudy (campbell_st):

same process use the point (0, -56)

OpenStudy (anonymous):

i see, thanks alot

OpenStudy (campbell_st):

P(x) = a(x + 5)(x-4)^2

OpenStudy (anonymous):

Find a degree 3 polynomial that has zeros - 4, 3 and 8 and in which the coefficient of x^2 is -14. How would i set this up? (x+4)(x-3)(x-8)=f(x) but how would i use the x^2=14?

OpenStudy (campbell_st):

ok... so you need to distribute your polynomial so you have \[P(x) = a(x^3 -7x^2 -20x +72)\] so you need \[-7a = 14\] solve for a

OpenStudy (anonymous):

where do you get the first equation? did you distribute the problem i wrote?

OpenStudy (anonymous):

or did you somehow incorperate the coeffient of x^2?

OpenStudy (campbell_st):

I distributed the factors from the polynomial you gave

OpenStudy (campbell_st):

oops... should be + 96

OpenStudy (campbell_st):

so then its a case of equating coefficients

OpenStudy (anonymous):

i see, so how is the coef. of x^2=14 relevant to the problem?

OpenStudy (campbell_st):

oops.. I thought it read 14x^2

OpenStudy (campbell_st):

the question says that the coefficient of x^2 is -14 so it means the term is -14x^2 and in the expanded polynomial I did x^2 has a coefficient of -7 so the value of a must be 2

OpenStudy (anonymous):

i see, but why did you choose 7? or was the what the original equation factored out to ?

OpenStudy (anonymous):

i see, so how is the coef. of x^2=14 relevant to the problem?

OpenStudy (campbell_st):

ok... your polynomial in factored form is P(x) =a[(x +4)(x-3)(x-8)] expand the factors.... and you'll have a cubic with 4 terms... only when you have expanded..or..distributed can you use the information about the coefficient of x^2 so the 1st part is \[P(x) =a[(x^2 + x - 12)(x - 8)]\] I hope this helps

OpenStudy (anonymous):

yes, i see now. thank you